in 5 percentage of completion KSON HAS THE FOLLOWING INFORMATION PERTAINING TO A LONG TERM PROJECT 10)COUNTED FOR UNDER THE P need help with solution please
Answer #1
Please give positive ratings so I can keep answering. Thanks!
Jackson 2015 2016
Calculation of Gross Profit Ans to 2a and 2b Journal Entries Debit $ Credit $ Debit $ Credit $
Particulars 2015 2016 2017 Costs Incurred CIP (Work-in-progress)       140,000      173,500
Contract Price              620,000             620,000             620,000 A Cash       140,000        173,500
Cost Incurred to date              140,000             313,500             418,000 B
Cost Incurred during the year              140,000             173,500             104,500 C
(Cost incurred to date of the current year less Cost incurred to date of the previous year) Period end adjustments for Revenue Construction Expense       140,000      173,500
Estimated cost to complete              260,000             104,500                        -   D (Gross Profit entry) CIP (Work-in-progress)         77,000        74,500
Estimated total cost              400,000             418,000             418,000 E=B+D Construction Revenue       217,000        248,000
Percentage complete 35.00% 75.00% 100.00% F=B/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized              217,000             248,000             155,000 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 2015: 35% completed. Revenue recognized = 35% x $ 620,000 = $ 217,000.
Year 2016: 75% completed. Revenue recognized = 75% x $ 620,000 – $ 217,000 (previously recognized) = $ 248,000.
Year 2017: 100% completed. Revenue recognized = 100% x $ 620,000 – $ 217,000- $ 248,000 (previously recognized) = $ 155,000.
Gross Profit/(Loss)                77,000               74,500               50,500 H=G-C